### v2

# Actively Learning Combinatorial Optimization Using a Membership Oracle

Abstract
We consider solving a combinatorial optimization problem with an unknown linear constraint using a membership oracle that, given a solution, determines whether it is feasible or infeasible with absolute certainty. The goal of the decision maker is to find the best possible solution subject to a budget on the number of oracle calls. Inspired by active learning based on Support Vector Machines (SVMs), we adapt a classical framework in order to solve the problem by learning and exploiting a surrogate linear constraint. The resulting new framework includes training a linear separator on the labeled points and selecting new points to be labeled, which is achieved by applying a sampling strategy and solving a 0-1 integer linear program. Following the active learning literature, one can consider using SVM as a linear classifier and the information-based sampling strategy known as Simple margin. We improve on both sides: we propose an alternative sampling strategy based on mixed-integer quadratic programming and a linear separation method inspired by an algorithm for convex optimization in the oracle model. We conduct experiments on the pure knapsack problem and on a college study plan problem from the literature to show how different linear separation methods and sampling strategies influence the quality of the results in terms of objective value.
1 Introduction
Mathematical programming is a powerful way of modeling and solving optimization problems. It has been proven to be effective for solving optimization problems with hard (e.g., combinatorial) constraints, in particular with help from off-the-shelf solvers (e.g., Gurobi [16]). The formalization of an optimization problem into a correct mathematical programming model is a process that requires expertise and knowledge about the specific application field. Lack of such resources or the intrinsic nature of the problem can make it difficult or impossible to obtain a fully specified mathematical model applying the traditional modeling approach, which relies extensively on human intervention. Data-driven techniques can help reduce this gap between applicability of mathematical programming and modeling capability, either by enabling solution techniques that do not require an explicit representation of objectives and constraints (e.g., [8, 25]) or by using machine learning methods to automate the formulation of mathematical programming models [11, 26].
In this paper, we take a step inside the development of specific techniques for combinatorial optimization in an interactive setting. We are interested in solving combinatorial optimization problems that are only partially described while learning a representation of the non-described portion. In order to limit the computational complexity of the optimization task, we assume a single unknown knapsack constraint. At the same time, we rely on a weak source of feedback consisting in receiving yes/no feasibility responses on single solutions. More formally, we want to solve:
$z^{*}=\max~{}$ $z(\boldsymbol{x}):=\boldsymbol{v}\cdot\boldsymbol{x}$ (1a)
s.t. $\sum_{i=1}^{d}\bar{w}_{i}\cdot x_{i}\leq 1,$ (1b)
$\boldsymbol{x}\in X\subseteq\{0,1\}^{d},$ (1c)
where weights $\bar{\boldsymbol{w}}\in W\subseteq[0,1]^{d}$ are unknown, the set $X$ encodes the known combinatorial structure of the problem in addition to the integrality constraints, while $W$ represents the prior knowledge at our disposal on the unknown weights $\bar{\boldsymbol{w}}$. We assume $W$ to be a polyhedron, i.e., $W=\{\boldsymbol{w}\in\mathbb{R}^{d}\ |\ \boldsymbol{a}\cdot\boldsymbol{w}\leq b\ \ \forall(\boldsymbol{a},b)\in\Delta\}\subseteq[0,1]^{d}$, where $|\Delta|<+\infty$. The unknown constraint (1b) is characterized by a membership oracle (with respect to $X$) that can determine the feasibility of any given solution $\boldsymbol{x}\in X$ with absolute certainty. Non-empty initial sets $S^{+}\subseteq X$ and $S^{-}\subseteq X$ are given, containing feasible and infeasible solutions to the unknown constraint (1b), respectively. We are also given a maximum number $N$ of oracle calls to respect, which reflects the fact that, when solving oracle-based problems in real contexts, the practical cost of a single call often plays a significant role.
Inspired by active learning methods for binary classification based on Support Vector Machines (SVMs) [22], we define an iterative solution paradigm prescribing some key steps. The first one is a linear separation step that consists in training a linear classifier to separate the known feasible from the known infeasible solutions and to find the equation of the separating hyperplane $h:\hat{\boldsymbol{w}}\cdot\boldsymbol{x}=1$, with $\hat{\boldsymbol{w}}\in W$. In a second phase, we select an unlabeled solution applying a suitable sampling strategy and solve the 0-1 integer programming (IP) model using the surrogate knapsack constraint $\hat{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1$ in place of the unknown one. Both the solution selected by the sampling strategy and the optimal solution of the surrogate IP model are submitted to the oracle to be labeled. We emphasize that, from a theoretical standpoint, the only way to have a guarantee of optimality on a specific solution would be enumerating an exponential number of solutions, so, practically, the goal of our algorithmic framework is to compute the best solution, say $\hat{x}$, within the maximum number of oracle calls.
We discuss an established approach in active learning that consists in using a SVM for linear classification together with a related sampling strategy that is function of the current separating hyperplane (a so-called information-based strategy). In particular, we employ a strategy known as simple margin [37], which applies to our combinatorial setting in a straightforward way. We propose an alternative sampling strategy called closest cutting plane that exploits the discrete nature of the problem and the structure of the so-called version space, i.e., the set of all surrogate weights $\hat{\boldsymbol{w}}\in W$ such that the hyperplane $h:\hat{\boldsymbol{w}}\cdot\boldsymbol{x}=1$ linearly separates the current labeled solutions. Subsequently, we propose an alternative linear separation method obtained as a variant of an algorithm for convex optimization with a separation oracle [8]. [^1] By separation in the separation oracle, we mean finding a hyperplane that separates a point from a convex set, which is different from separation in the linear separation of labeled points. Given a convex set $K\subseteq\mathbb{R}^{d}$, a separation oracle for the set takes in input any $\boldsymbol{w}\in\mathbb{R}^{d}$ and outputs either “yes” if $\boldsymbol{w}\in K$ or a linear inequality valid for $K$ but violated by $\boldsymbol{w}$. Our version of the algorithm computes the vector of surrogate weights relying exclusively on the membership oracle at our disposal, which is used to approximate a separation oracle.
We conduct experiments to assess the effectiveness of different combinations of linear separation method and sampling strategy to approximate an optimal binary solution. For this purpose, we consider two specific realizations of the original problem (1). The first one is the pure knapsack case, in which $X=\{0,1\}^{d}$. The second one is inspired by the generation of personalized learning systems with uncertain item weights [1] and features known constraints in addition to the unknown knapsack constraint.
The remainder of the paper is organized as follows. In Section 2, we give some background and we review the related literature. In Section 3, we give the details of our solution approaches. In Section 4, we present our experimental setting and some empirical results. Finally, we draw our conclusions in Section 5.
2 Background and Literature Review
In this Section we review the literature and give some background about convex optimization in the oracle setting, then we briefly present some principles and methodologies of active learning on which our algorithmic contributions rely heavily.
2.1 Oracle-Based Optimization
Oracle-based optimization has received substantial attention over the years for its role in applications where an explicit description of the feasible region is not available or difficult to formalize. A well-studied setting consists in minimizing a convex function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ over a convex set $K$ equipped with a separation oracle [3, 7, 8, 13, 14, 21, 25, 29, 39], which is proven to be solvable with a pseudo-linear number of oracle calls with respect to the dimension $d$. Recent advances include convex optimization with a membership oracle of $K$ and an evaluation oracle of $f$ [24]. In the case of convex sets, separation and membership oracles are equivalent from a complexity standpoint. In fact, according to the classical result in [15], there exists a polynomial time equivalence between optimization, separation and membership for convex bodies. Since in this paper we are interested in the combinatorial optimization scenario, convenient theoretical results like those described above are not available due to the non-convexity of the feasible region. Actually, as shown in the example of Appendix A, any randomized algorithm requires $\Omega(n^{1/\epsilon})$ calls to the membership oracle in expectation to find a $(1-\epsilon)$ optimal solution of a knapsack problem with an unknown knapsack constraint in the worst case.
2.2 Active Learning for Linear Separation
The solution approaches that we propose in this paper to solve the general problem introduced in Section 1 are based on learning a surrogate linear constraint or, in other terms, an approximation of the halfspace containing the feasible points according to the original unknown constraint. Halfspace learning is a research area with a long tradition that traces back to the Perceptron algorithm [32]. It has received extensive interest over the years (see, e.g., [2, 4, 5, 9, 10, 17, 35]), especially regarding popular algorithms like, e.g., Support Vector Machines (SVMs) [6, 27, 40]. A significant part of the work in halfspace learning has been done in the context of active learning [12, 18, 23, 34, 36]. Compared to traditional supervised learning techniques, that can usually require massive amounts of labeled data to produce models with the desired generalization capabilities, active learning algorithms employ relatively small initial labeled datasets and are based on querying an oracle on new unlabeled data in order to make the learning process less data consuming [36]. A particular sub-category of active learning methods is represented by algorithms whose distinctive characteristic is to rely on a fixed dataset of unlabeled data and to implement a policy, called sampling strategy, to decide what unlabeled available points submit to the oracle for labeling, in a sequential fashion. An established active learning method for binary classification consists in alternating the training of a binary classifier on the dataset of already labeled points and a sampling step. The classification task is traditionally accomplished using a SVM [19, 33, 38, 41], while several related sampling strategies have been proposed in the literature (see, e.g., [22] and [36]). Existing strategies fall in one of three categories: information-based strategies, that sample the most informative points, for example selecting those whose label is the most uncertain; representation-based strategies, that try to diversify the areas where the points are sampled; and finally, strategies that combine the two previous approaches.
In the presence of separable data, the set of all the classifiers in the hypothesis space that correctly separate the current observations is called version space [28, 36]. Training a classifier on the current labeled points has the “dual” meaning of finding a point inside the current version space. Sampling an unlabeled point and acquiring its label from the oracle corresponds to constraint and possibly reduce the version space in size. The goal behind the idea of alternating the training of a binary classifier and a sampling strategy is to find a convenient “dual” point inside the version space and subsequently to sample a point in the original “primal” space whose effect is to significantly shrink the version space.
3 Active Learn-and-Optimize
As mentioned in Section 2, a typical solution scheme in active learning for binary classification consists in alternating the training of a binary classifier and the application of a sampling strategy. We adapt this methodology adding an optimization step and we formalize the resulting framework for combinatorial optimization with knapsack constraint learning using a membership oracle. The framework, that we call active learn-and-optimize, is presented in Algorithm 1. The pseudo-code employs the following notations:
  1. 1.
    $\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$ denotes the linear separation algorithm of choice, that outputs $\hat{\boldsymbol{w}}_{t}\in W$ such that $\hat{\boldsymbol{w}}_{t}\cdot\boldsymbol{s}^{+}\leq 1$ for every $\boldsymbol{s}^{+}\in S^{+}_{t}$ and $\hat{\boldsymbol{w}}_{t}\cdot\boldsymbol{s}^{-}\geq 1$ for every $\boldsymbol{s}^{-}\in S^{-}_{t}$;
  2. 2.
    $\mathcal{T}(X,S^{+}_{t},S^{-}_{t})$ corresponds to the chosen sampling strategy, that returns a point $\boldsymbol{u}_{t}$ from $X\setminus(S^{+}_{t}\cup S^{-}_{t})$;
  3. 3.
    $\Omega(\boldsymbol{p})$ indicates a query on point $\boldsymbol{p}$ that yields its label $\ell_{\boldsymbol{p}}\in\{1,-1\}$, where 1 corresponds to feasible and $-1$ to infeasible.
Algorithm 1 Active learn-and-optimize
1:Input: maximum number of queries $N$, initial sets $S^{+}_{0}=S^{+}$ and $S^{-}_{0}=S^{-}$ of labeled points, surrogate weight domain $W$, rewards $\boldsymbol{v}$
2:$t\leftarrow 0$
3:$n_{q}\leftarrow 0$
4:while $n_{q}<N$ do
5:     $\hat{\boldsymbol{w}}_{t}\leftarrow\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$triangleright Linear separation
6:     $t\leftarrow t+1$
7:     $P^{+}_{t}\leftarrow\emptyset$
8:     $P^{-}_{t}\leftarrow\emptyset$
9:     for $\theta\in\{\text{sampling},\text{optimization}\}$ do
10:         if $\theta=\text{sampling}$ then
11:              $\boldsymbol{u}_{t}\leftarrow\mathcal{T}(X,S^{+}_{t-1},S^{-}_{t-1})$triangleright Sampling strategy
12:         else
13:              Get $\boldsymbol{u}_{t}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in X\setminus(S^{+}_{t-1}\cup S^{-}_{t-1})\text{,}\ \hat{\boldsymbol{w}}_{t}\cdot\boldsymbol{x}\leq 1\right\}$triangleright Optimization
14:         $\ell_{\boldsymbol{u}_{t}}\leftarrow\Omega(\boldsymbol{u}_{t})$triangleright Oracle call
15:         $n_{q}\leftarrow n_{q}+1$
16:         if $\ell_{\boldsymbol{u}_{t}}=1$ then
17:              $P^{+}_{t}\leftarrow P^{+}_{t}\cup\{\boldsymbol{u}_{t}\}$
18:         else
19:              $P^{-}_{t}\leftarrow P^{-}_{t}\cup\{\boldsymbol{u}_{t}\}$
20:         if $n_{q}=N$ then $\triangleright$ Stopping criterion
21:              break
22:     $S^{+}_{t}\leftarrow S^{+}_{t-1}\cup P^{+}_{t}$
23:     $S^{-}_{t}\leftarrow S^{-}_{t-1}\cup P^{-}_{t}$
24:Get $\hat{\boldsymbol{x}}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in X,\boldsymbol{x}\leq\boldsymbol{y}\text{ for some }\boldsymbol{y}\in S^{+}_{t}\right\}$triangleright Final solution
25:$\hat{\boldsymbol{w}}\leftarrow\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$triangleright Final linear separation
26:return $\hat{\boldsymbol{x}}$, $\hat{\boldsymbol{w}}$
We can conceptually identify a few main steps inside the framework: the linear separation step (line 5) finds the surrogates weights by computing $\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$; the sampling step (line 11) is the application of the chosen sampling strategy, i.e., $\mathcal{T}(X,S^{+}_{t},S^{-}_{t})$, to select an unlabeled point; the optimization step (line 13) finds an optimal solution of the partially specified 0-1 IP model using the current surrogate constraint in place of the unknown knapsack one. Both the sampling and the optimization steps are followed by an oracle call that returns the label of the newly found point (line 14). The stopping criterion (line 20) is triggered by reaching the maximum number of queries. In cases when $\boldsymbol{v}\geq 0$, one can simplify the final solution step (line 24) to $\hat{\boldsymbol{x}}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in S^{+}_{t}\right\}$. We remark that, since the sampling step and the optimization step always return new points with respect to the previous iterations, and since the set $X$ is finite, any instance of the framework ensures finite convergence to the optimal solution when permitting an infinite number of oracle calls. Finally, we observe that, in principle, the framework admits modifications such as employing multiple sampling or optimization steps between any two consecutive linear separation steps, or selecting just one single point to query at every time step $t$.
As is common in active learning, SVM can be used to accomplish the linear separation task. Specifically, we may solve the following linear programming model and set $\boldsymbol{w}/b$ as the surrogate weights:
min $\ \sum_{i=1}^{d}w_{i}^{2}$
s.t. $\ b-\sum_{i=1}^{d}w_{i}\cdot s^{+}_{i}\geq 1,$ $\boldsymbol{s}^{+}\in S^{+}_{t},$
$\ b-\sum_{i=1}^{d}w_{i}\cdot s^{-}_{i}\leq-1,$ $\boldsymbol{s}^{-}\in S^{-}_{t},$
$\ \boldsymbol{w}\in\mathbb{R}^{d},~{}b\in\mathbb{R}.$
For the sampling step, one can consider multiple existing approaches. However, established SVM-related strategies like, e.g., MaxMin margin or Ratio margin, require to train a linear classifier for every unlabeled point [22]. In general, most of the sampling strategies present in the literature are incompatible with a combinatorial optimization setting, in which there is an exponential number of unlabeled examples to choose from. To the best of our knowledge, the only existing strategy that can be efficiently adapted to such a setting is Simple margin (SIM) [37]. It consists in selecting the point that maximizes the uncertainty with respect to the classification provided by the current SVM, i.e., the unlabeled point closest to the decision boundary. The task can be accomplished by solving the following model:
min $\ \left|1-\sum_{i=1}^{d}\hat{w}_{t,i}\cdot u_{i}\right|$ (2a)
s.t. $\ \sum_{i=1}^{d}u_{i}\cdot(1-s^{+}_{i})\geq 1,$ $\boldsymbol{s}^{+}\in S^{+}_{t},$ (2b)
$\ \sum_{i=1}^{d}u_{i}\cdot s^{-}_{i}\leq\sum_{i=1}^{d}s^{-}_{i}-1,$ $\boldsymbol{s}^{-}\in S^{-}_{t},$ (2c)
$\boldsymbol{u}\in X.$ (2d)
The objective function (2a) minimizes the distance between the generic point $\boldsymbol{u}\in X$ and the hyperplane identified by $\hat{\boldsymbol{w}}$. No-good cuts (2b) and (2c) are used to prevent finding a solution whose feasibility or infeasibility is implied by labeled points in the sets $S^{+}_{t}$ and $S^{-}_{t}$. We remark that, although integrality constraints make the problem NP-hard in the general case, mixed-integer optimization models can be often solved efficiently, e.g., by off-the-shelf discrete optimization solvers.
3.1 Closest Cutting Plane Sampling Strategy
For the pair of sets $S^{+}_{t}$ and $S^{-}_{t}$ of feasible and infeasible points at time step $t$, the current version space is defined as the set:
$V_{t}$ =$=\{\boldsymbol{w}\in W\ |\ \boldsymbol{w}\cdot\boldsymbol{s}^{+}\leq 1\ \text{for all}\ \boldsymbol{s}^{+}\in S^{+}_{t}\text{,}\ \boldsymbol{w}\cdot\boldsymbol{s}^{-}>1\ \text{for all}\ \boldsymbol{s}^{-}\in S^{-}_{t}\}\text{.}$
Let us also define $S^{+}_{\infty}=\{\boldsymbol{x}\in X\ |\ \bar{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1\}$ and $S^{-}_{\infty}=\{\boldsymbol{x}\in X\ |\ \bar{\boldsymbol{w}}\cdot\boldsymbol{x}>1\}$, as the sets of all feasible and infeasible points, respectively. Every time we update $S^{+}_{t}$ and $S^{-}_{t}$ with new labeled points to obtain $S^{+}_{t+1}$ and $S^{-}_{t+1}$, the next version space $V_{t+1}$ will be defined by additional inequalities with respect to $V_{t}$. The sequence of version spaces converges (in a finite but potentially exponential number of steps) to the set of surrogate weights whose corresponding surrogate constraint classifies every point in $X$ on the same side as the unknown knapsack constraint $\bar{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1$. Namely,
$V_{\infty}$ =$=\{\boldsymbol{w}\in W\ |\ \boldsymbol{w}\cdot\boldsymbol{x}\leq 1\ \text{if and only if}\ \bar{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1\ \text{for all}\ \boldsymbol{x}\in X\}$
=$=\{\boldsymbol{w}\in W\ |\ \hat{\boldsymbol{w}}\cdot\boldsymbol{s}^{+}\leq 1\ \text{for all}\ \boldsymbol{s}^{+}\in S^{+}_{\infty}\text{,}\ \boldsymbol{w}\cdot\boldsymbol{s}^{-}>1\ \text{for all}\ \boldsymbol{s}^{-}\in S^{-}_{\infty}\}\text{.}$ (3)
Intuitively, if the vector of surrogate constraints $\hat{\boldsymbol{w}}$ is sufficiently close to $V_{\infty}$, then sampling an unlabeled point $\boldsymbol{u}$ from $X$ whose corresponding cut is close to $\hat{\boldsymbol{w}}$ can have the effect of significantly reducing the size of the version space, hence accelerating the convergence. Following this intuition, we propose a sampling strategy alternative to SIM, named Closest cutting plane (CUT). Given $\hat{\boldsymbol{w}}$ in the closure $\bar{V}_{t}$ of the current version space, CUT finds an unlabeled point $\boldsymbol{u}\in X$ whose corresponding hyperplane $h_{\boldsymbol{u}}=\{\boldsymbol{w}\in\mathbb{R}^{d}\ |\ \boldsymbol{w}\cdot\boldsymbol{u}=1\}$ has minimum Euclidean distance from $\hat{\boldsymbol{w}}$. To accomplish this task, we solve the following mixed-integer quadratic program:
min $\ \sum_{i=1}^{d}(p_{i}-\hat{w}_{t,i})^{2}$ (4a)
s.t. $\ \sum_{i=1}^{d}p_{i}\cdot u_{i}=1,$ (4b)
$\ \sum_{i=1}^{d}u_{i}\cdot(1-s^{+}_{i})\geq 1,$ $\boldsymbol{s}^{+}\in S^{+}_{t},$ (4c)
$\ \sum_{i=1}^{d}u_{i}\cdot s^{-}_{i}\leq\sum_{i=1}^{d}s^{-}_{i}-1,$ $\boldsymbol{s}^{-}\in S^{-}_{t},$ (4d)
$\ \boldsymbol{u}\in X,$ (4e)
$\ \boldsymbol{p}\in\mathbb{R}^{d}.$ (4f)
The distance between $\hat{\boldsymbol{w}}$ and the hyperplane $h_{\boldsymbol{u}}$ is modeled as the distance between $\hat{\boldsymbol{w}}$ and its projection $\boldsymbol{p}$ on $h_{\boldsymbol{u}}$, which is minimized by the objective function 4a. The quadratic constraint (4b) ensures that $\boldsymbol{p}$ belongs to $h_{\boldsymbol{u}}$. No-good cuts (4c) and (4d) are used exactly as in SIM.
3.2 Approximation of a Separation Oracle and the Resulting Linear Separation Algorithm
In principle, given any vector of surrogate weights in $V_{\infty}$, one can solve the original combinatorial optimization problem (1) by replacing the unknown constraint with the surrogate constraint. Therefore, in order to solve (1), one can try to find a point in the convex set $V_{\infty}$ using a membership oracle of the constraint $\bar{w}\cdot\boldsymbol{x}\leq 1$ with respect to $X$. It is well-known that one can efficiently find a point in a well-bounded convex body $K$ with a separation oracle. Specifically, assuming that a compact convex set $K\subseteq\mathbb{R}^{d}$ is contained in the Euclidean ball of radius $R$ centered at the origin, there exist algorithms using $O(d\log(dR/\epsilon))$ calls to the separation oracle to either find a point in $K$ or prove that $K$ does not contain any Euclidean ball of radius $\epsilon$ [8, 20, 25]. However, such algorithms are not directly applicable in our case because we do not have direct access to a separation oracle of $V_{\infty}$. On the other hand, given a weak separation oracle of a convex body $K$ (that finds an almost violated linear inequality valid for $K$), there exist algorithms that can find a point almost in $K$ in oracle-polynomial time [15]. To hopefully find a point in $V_{\infty}$ (or at least get a good approximation of $V_{\infty}$), we propose to construct an approximate separation oracle of $V_{\infty}$ using the membership oracle of (1b), and replace the separation oracle in existing convex optimization algorithms by the constructed approximate separation oracle. In particular, we adapt the algorithm proposed in [8] as our choice of the convex optimization algorithm in the oracle model. Implementation details are described below.
The original version of the algorithm presented in [8] is intended to minimize a convex function $f$ over a full-dimensional convex body $K\subseteq\mathbb{R}^{d}$. Specifically, it generates a sequence of points in $\mathbb{R}^{d}$ that converges to a minimum inside $K$, under the assumption that $K$ is contained inside a ball of radius $R$ centered in the origin. Such a set $K$ is given implicitly by means of a separation oracle, which for every point $\hat{\boldsymbol{w}}\in\mathbb{R}^{d}$, either states that $\hat{\boldsymbol{w}}\in K$ or returns an inequality $\boldsymbol{\alpha}\cdot\boldsymbol{w}\leq\beta$ valid for $K$ but violated by $\hat{\boldsymbol{w}}$. To address the solution of problem 1, we assume $f$ to be constant and we let $K$ be the set $V_{\infty}$ defined in the previous Section 3.1. Since we do not have access to a separation oracle for $K$, we approximate it using the membership oracle at our disposal, that characterizes the set $S^{+}_{\infty}$ of all the feasible solutions. For each feasible point $\boldsymbol{s}^{+}$, the inequality $\hat{\boldsymbol{w}}\cdot\boldsymbol{s}^{+}\leq 1$ is valid for $K$. Similarly, for each infeasible point $\boldsymbol{s}^{-}$, the inequality $-\hat{\boldsymbol{w}}\cdot\boldsymbol{s}^{-}\leq-1$ is also valid for $K$. For every pair of sets $S^{+}_{t}$ and $S^{-}_{t}$ of labeled points, the valid inequalities corresponding to such points in addition to the inequalities that define $W$ (i.e., $\boldsymbol{a}\cdot\boldsymbol{w}\leq b$ for $(\boldsymbol{a},b)\in\Delta$) determine the current version space $V_{t}\supseteq K$ as defined in Section 3.1.
[…]
Algorithm 2 Alternative linear separator (SEP)
1:Input: $W=\{\boldsymbol{w}\ |\ \boldsymbol{a}\cdot\boldsymbol{w}\leq b\text{,}\ (\boldsymbol{a},b)\in\Delta\}$, iteration counter $t$, sets $S^{+}_{t}$ and $S^{-}_{t}$ of feasible and infeasible points respectively
2:$\Lambda\leftarrow\{(\boldsymbol{a},b)/\|(\boldsymbol{a},b)\|_{*}\ |\ (\boldsymbol{a},b)\in\Delta\}$
3:for $\boldsymbol{s}^{+}\in S^{+}_{t}$ do
4:     $\Lambda\leftarrow\Lambda\cup\{(\boldsymbol{s}^{+},1)/\|(\boldsymbol{s}^{+},1)\|_{*}\}$
5:for $\boldsymbol{s}^{-}\in S^{-}_{t}$ do
6:     $\Lambda\leftarrow\Lambda\cup\{(-\boldsymbol{s}^{-},-1)/\|(-\boldsymbol{s}^{-},-1)\|_{*}\}$
7:Get a point $\boldsymbol{\gamma}_{t}\in\arg\min\{\Phi(\gamma)\ |\ \gamma\in\text{conv}(\Lambda)\}$
8:$\hat{\boldsymbol{w}}_{t}\leftarrow-\left(\nabla\Phi(\gamma_{t})_{1},\ldots,\nabla\Phi(\gamma_{t})_{d}\right)/\nabla\Phi(\gamma_{t})_{d+1}$
9:return $\hat{\boldsymbol{w}}_{t}$
The original algorithm from [8] guarantees a convergence rate towards a vector in $K$. However, in our case, no guarantees can be given regarding the convergence of $\hat{\boldsymbol{w}}_{t}$ towards a vector in $K=V_{\infty}$, because we are approximating the separation oracle required by the original algorithm using a membership oracle. We base the evaluation of the adapted algorithm on the experimental results in Section 4.
4 Experimental Results
In this section, we evaluate different combinations of linear separator methods and sampling strategies in the application of Algorithm 1. In particular, we combine SVM with SIM as traditionally done in active learning, then we apply SVM with our sampling strategy CUT and finally we test our alternative linear separator SEP with CUT. We evaluate such three combinations on two test problems, namely the pure knapsack problem, in which $X=\{0,1\}^{d}$, and the college study plan problem. We show how differences arise between the combinations in terms of the objective value of the best found feasible solutions.
4.1 College Study Plan Problem
The college study plan problem (CSPP) is a 0-1 combinatorial optimization problem in which a prospective college student needs to design their study plan in order to maximize the number of credits acquired respecting an estimated time budget. Each course yields a predefined number of credits and requires an effort in terms of study hours given by a certain fixed number and an amount of hours for independent study that depends on the particular student. Given a possible study plan, the student needs to decide whether or not taking the courses specified in the plan is feasible with respect to the time budget. We assume that an unknown knapsack constraint models the satisfaction of the time budget. Additional known constraints model the dependencies between the available courses present in the course catalog. Any course $i$ can be included in a study plan only if at least one course from each of a sequence of prerequisite sets $P^{i}_{1}$, $P^{i}_{2}$, $\ldots$, $P^{i}_{N^{i}}$ is included in the plan. Every course $i$ also has a set $C^{i}$ of corequisites (courses that have to be taken if and only if course $i$ is taken) and a set $A^{i}$ of alternatives (courses that must be taken if $i$ is taken). Reasonable assumptions avoid inconsistency of the dependencies, namely: (a) $\bigcup_{k=1}^{N^{i}}P^{i}_{k}$, $C^{i}$ and $A^{i}$ are pairwise disjoint sets for every $i\in\{1,\ldots,d\}$; (b) if $i\in\bigcup_{k=1}^{N^{j}}P^{j}_{k}$ then $j\notin\bigcup_{k=1}^{N^{i}}P^{i}_{k}$, for every $i,j\in\{1,\ldots,d\}$; (c) $i\in C^{j}$ if and only if $j\in C^{i}$ for every $i,j\in\{1,\ldots,d\}$; (d) $i\in A^{j}$ if and only if $j\in A^{i}$ for every $i,j\in\{1,\ldots,d\}$. The goal is to solve the following 0-1 integer linear programming model with an unknown knapsack constraint:
max $\sum_{i=1}^{d}v_{i}\cdot x_{i}$
s.t. $\sum_{i=1}^{d}\bar{w}_{i}\cdot x_{i}\leq 1,$
$x_{i}\leq\sum_{j\in P^{i}_{k}}x_{j},$ $k\in\{1,\ldots,N^{i}\},$
$x_{i}=x_{j},$ $j\in C^{i},$
$x_{i}+x_{j}\leq 1,$ $j\in A^{i},$
$x_{i}\in\{0,1\},$ $i\in\{1,\ldots,d\}.$
Here, each reward $v_{i}\geq 0$ represents the number of credits associated to course $i$. Each weight in the knapsack constraint can be modeled as $\bar{w}_{i}=\min\{\eta_{i}+\xi_{i},1\}$, where $\eta_{i}\geq 0$ represents the known number of course hours $i$ and $\xi_{i}\geq 0$ the unknown student-dependent component, assuming that both are normalized by dividing them by the unknown total time budget.
4.2 Computational Results
The query limit was set equal to 2000 for all the experiments. Other experimental setups are presented in Appendix C. The algorithms were tested on 3 sets of instances. We briefly describe their structure and we refer to Appendix B for the generation details. First, we obtained a set of knapsack instances with $d=100$ items and 3 types of rewards and weights: uncorrelated (KNAP-U), weakly correlated (KNAP-W) and strongly correlated (KNAP-S). For each type, we produced 30 instances. For CSPP we generated two sets of instances of dimension $d=150$. The first one (MIT-Math) contains 30 instances structured according to the list of undergraduate courses in Mathematics at the Massachusetts Institute of Technology (MIT). The second one (SYNTH-MIT) is composed by additional 30 synthetic instances, with randomly generated structures inspired by the dependencies of the previous MIT courses.
For different sets of instances, we summarize in Figures 1-5 how the relative objective error, averaged over the specific instances, decreases as the number of used oracle calls increases. Here the relative objective error is computed by $(z^{*}-\boldsymbol{v}\cdot\hat{\boldsymbol{x}})/z^{*}\cdot 100\%$ with $\hat{\boldsymbol{x}}$ denoting the best feasible solution found by individual methods. We first note that the average performance of all test methods on randomly generated instances is significantly better than the worst case lower bound presented in Example 1. Even though the dimension of the test problems is at least 100, all test methods can efficiently find feasible solutions with average relative objective error lower than 10% using up to a few hundred oracle calls. In particular, the relative objective error usually drops quickly at the very beginning and then tails off. Unlike traditional knapsack problems [31], strongly correlated knapsack instances seem easier, in terms of the number of oracle queries needed, to find almost optimal solutions with small (e.g., 1%) relative objective error, than uncorrelated and weakly correlated knapsack instances. The overall performance of our newly proposed methods is very satisfactory. On the knapsack instances, the method SEP+CUT outperforms the other alternatives (and SVM+CUT outperforms SVM+SIM) in relative objective error when the 2000 query limit is reached. On the CSPP instances, SEP+CUT converges to 1% relative objective error much faster than the other alternatives even though SVM+SIM attains a smaller average ending relative objective gap on MIT-Math CSPP instances. Additional experimental results are presented in Appendix D.
Figure 1: Average relative objective error convergence for KNAP-U instances.
Figure 2: Average relative objective error convergence for KNAP-W instances.
Figure 3: Average relative objective error convergence for KNAP-S instances.
Figure 4: Average relative objective error convergence for MIT-Math CSPP instances.
Figure 5: Average relative objective error convergence for synthetic CSPP instances.
5 Conclusions
We proposed an active learn-and-optimize framework and its alternative realizations, for solving combinatorial optimization problems with an unknown linear constraint accessed by a membership oracle. Extensive numerical experiments are conducted to test the effectiveness of our proposed approaches. Even though such problems can be very hard in the worst case, we find that “average" combinatorial instances are not much harder than its convex counterpart. An interesting open question is how to balance the frequency of sampling and optimization steps. It would also be interesting to see if there exists any algorithm that matches the lower bound result in Example 1.
Appendix A Hardness Example
Example 1.
For any fixed $\epsilon$ such that $1/\epsilon$ is an integer constant, consider a knapsack problem instance with $n$ items, where $1/\epsilon$ items have weights equal to $1$, $(n-1/\epsilon)$ items have weights equal to $1/(1-\epsilon)$, and all items have identical profit equal to $1$. Assume that the knapsack has capacity $1/\epsilon$, and the item weights are unknown such that the knapsack constraint can only be accessed through a membership oracle. This knapsack instance has a unique optimal solution by picking all the weight $1$ items, which attains the optimal objective value $1/\epsilon$. Note that all other feasible solutions pick at most $11/\epsilon-$ 1 items and all solutions that pick at most $11/\epsilon-$ 1 items are feasible. Therefore, one cannot distinguish the weight $1$ items from the weight $1/(1-\epsilon)$ items from the oracle calls until one calls the membership oracle to query the feasibility of the optimal solution that picks all $1/\epsilon$ weight 1 items, in the sense that any solution that picks at most $11/\epsilon-$ 1 item is feasible while any solution that picks at least $1/\epsilon$ items is infeasible except the optimal solution. To find any better-than-$(1-\epsilon)$ solution (i.e., to find the optimal solution), any algorithm would in expectation (with respect to a random permutation of the order of the items) have to query the feasibility of almost all solutions that pick exactly $1/\epsilon$ items, which means at least $\Omega\left({{n}\choose{1/\epsilon}}\right)=\Omega(n^{1/\epsilon})$ calls to the membership oracle of the knapsack constraint.
Appendix B Instance Generation
We applied a Python re-implementation of a well known knapsack generator [30] to obtain test instances for the pure knapsack problem (KNAP) with 100 items, with rewards and weights of three types: uncorrelated (KNAP-U), weakly correlated (KNAP-W) and strongly correlated (KNAP-S). For each type, we generated 30 instances with items having integer weights in the range $[1,10000]$ and varying capacities, then for each instance we normalized the weights by dividing them by the corresponding capacity.
For CSPP, we built two sets of instances inspired by the experimental setting of [1]. The first set of instances (MIT-Math) is designed according to the structure of the undergraduate courses in Mathematics at the Massachusetts Institute of Technology (MIT), whose list can be found here: http://student.mit.edu/catalog/m18a.html. Differently from [1], we took into account the entire list and we also considered every dependency inside other areas (Physics, Computer Science, Biology, etc.). For every course, the credits obtainable for passing the corresponding exam are specified as a triple $(a_{i},b_{i},c_{i})$, where the three values represent credits assigned for lectures and recitations, for laboratory, design, or field work, and for outside (independent) preparation, respectively. Each credit corresponds to 14 hours of activity, but while this equivalence is exact for $a_{i}$ and $b_{i}$, the hours actually dedicated to outside preparation vary from student to student. We operated few adjustments to the original data, mainly consisting in merging homologous courses, and we obtained a dataset with 150 courses whose structure of dependencies respect the definition of the college study plan problem. We generated 30 test instances, where the parameters related to each course $i$ are computed according to the following specifications from [1] ($B$ denotes the total hours budget equal to 1700, while $\mathcal{N}_{\text{T}}$ denotes the truncated normal distribution supported on $[0,B]$):
  1. (a)
    $\eta_{i}=14\cdot(a_{i}+b_{i})/B$;
  2. (b)
    $\xi_{i}\sim\mathcal{N}_{\text{T}}(\mu_{i},\sigma_{i}^{2})/B$, where $\mu_{i}=14\cdot c_{i}$ and $\sigma_{i}\sim U[1,14×2]$;
  3. (c)
    $v_{i}=a_{i}+b_{i}+c_{i}$.
The second set of instances (SYNTH-MIT) for CSPP is synthetic and mimics the structure of the prerequisites of the MIT courses obtained previously (without considering corequisites and alternatives). We give a brief description of the generation process. First of all, we divided the courses in pairwise disjoint groups such that for all the courses in the same group, their prerequisites span an identical set of other groups (here, by prerequisites of a course $i$, we mean all the courses in their prerequisite sets $P^{i}_{1}$, $\ldots$, $P^{i}_{N^{i}}$). So, for example, group $C_{0}$ is composed by the courses that do not have prerequisites, group $S(C_{0})$ by the courses that only have prerequisites in $C_{0}$, group $S(C_{0},S(C_{0}))$ by the courses that have at least one prerequisite in $C_{0}$ and at least one in $S(C_{0})$, and so on recursively. The groups with more than 3 courses, which are 14 over 48, account for about 75% of the total number of courses. Using a number $d=150$ of courses, we generated our instances for CSPP at random such that $\lfloor 0.75\cdot d\rfloor$ of the courses form a structure of groups similar to the one of those 14 groups (both in terms of proportion between their sizes and in terms of prerequisite dependencies) and the remaining courses were assigned building additional random groups. We generated 30 instances using the following adaptation of criteria for synthetic instances applied in [1] (where $B=250\cdot d$):
  1. (a)
    $\eta_{i}\sim U[1,1000]/B$;
  2. (b)
    $\xi_{i}=\mathcal{N}_{\text{T}}(\mu_{i},\sigma_{i}^{2})/B$, where $\mu_{i}=\mathcal{N}(250,200^{2})$ and $\sigma_{i}\sim U[1,200]$;
  3. (c)
    $v_{i}=\bar{w}_{i}+200$.
Appendix C Experimental Setups
We implemented the algorithms described in Section 3 using Python 3.11.4. The optimization models were solved using Gurobi 11.0.0 with a limit of 4 threads on the single optimization. A small portion of instances of the models for SIM and CUT could require an excessive amount of time to be solved optimally, therefore we applied a node limit equal to 50000 for the Gurobi branch-and-bound algorithm. The experiments were conducted on a machine with 2 CPUs Intel(R) Xeon(R) Gold 6258R 2.70GHz 28 cores, 512 GB RAM and OS AlmaLinux 9.3.
Appendix D Additional Computational Results
We summarize in Table 1 different statistics obtained from the numerical experiments. In column “# Solved”, we report the number of instances where the exact optimal solution is found by the algorithm. In column “Ending Error (%)”, we report the relative objective error obtained when reaching the query limit. In column “Comput. Time (s)”, we report the total computational time required to solve the optimization subproblems in Algorithm 1. In column “# Queries”, we report the total number of queries. We use $\geq$" to indicate that more computational time and queries are needed to find the exact optimal solution of the instances. We observe that SEP+CUT solves the most instances in general, except for MIT-Math CSPP instances. The optimization subproblems required by the sampling strategy CUT seem more computationally demanding than the sampling strategy SIM.
Table 1: Overview of experimental results.
Instances Method # Solved Ending Error (%) Comput. Time (s) # Queries
KNAP-U SEP+CUT 30/30 0.00 2512 638
SVM+CUT 20/30 0.03 $\geq$11999 $\geq$1231
SVM+SIM 20/30 0.04 $\geq$01459 $\geq$1200
KNAP-W SEP+CUT 29/30 <0.01 $\geq$12425 $\geq$1240
SVM+CUT 28/30 <0.01 $\geq$10267 $\geq$1351
SVM+SIM 26/30 0.02 $\geq$02160 $\geq$1424
KNAP-S SEP+CUT 00/30 0.05 $\geq$30690 $\geq$2000
SVM+CUT 00/30 0.06 $\geq$25948 $\geq$2000
SVM+SIM 00/30 0.08 $\geq$03616 $\geq$2000
MIT-Math CSPP SEP+CUT 22/30 0.46 $\geq$06019 $\geq$1162
SVM+CUT 02/30 4.13 $\geq$12710 $\geq$1982
SVM+SIM 29/30 0.07 $\geq$03272 $\geq$1529
Synthetic CSPP SEP+CUT 04/30 0.05 $\geq$18499 $\geq$1976
SVM+CUT 02/30 0.33 $\geq$16786 $\geq$1984
SVM+SIM 02/30 0.54 $\geq$05169 $\geq$1985
At each row, statistics averaged over the specific set of instances are reported.